Assume that a customer shops at a local grocery store spending an average of $200 a week, resulting a retailer profit of $10 each week from this customer. Assuming the shopper visits the store all 52 weeks of the year, calculate the customer lifetime value if this shopper remains loyal over a 10-year lifespan. Also assume a 5 percent annual interest rate and no initial cost to acquire the customer.

Describe ways marketers can increase the lifetime value of a customer.
Answer #1

The average spending made in a week = $200 The customer spending by the customer during each year=200*52=10400 (one year = 52The results obtained are as follows: 0.05 Inflow 520 1 Rate of interest 2 Cash Flow 3 Period 1 4. Period 2 5 Period 3 6 Perio